Daily Events on 2008/12/09 (Tue)

last updated at  2008/12/08 19:29:04 JST


<< prev day

@

next day >>


Station

Pass No.
20081209-

Contact time (UT)

Max. Elev.
 (deg)

Events

AOS

LOS

USC20

0200

09:22:51

09:35:13

30.2

OP/OG upload
XRT: TI-CMD
EIS: TI-CMD
SOT: TI-CMD
DR playback (by OP)

KRN1

6010

09:40:58

09:52:44

16.7

OP/OG upload, XRT:TI-CMD, SOT:TI-CMD, EIS:TI-CMD (backup)
AOCS: Upload pointing table
AOCS: Upload orbital element
EIS: Table Upload

SNT1

4000

10:12:36

10:26:23

80.4

SOT: Table Upload
XRT: Table Upload

USC20

0400

21:24:21

21:38:08

84.5

DR playback (by OP)

Flare list(LM) : http://www.lmsal.com/solarsoft/last_events/


DR information

Remaining amount
(prediction)

- Mbits @ OP_START ( )


SOT

XRT

EIS

DR allocation [%]

70

15

15

DR allocation [Mbits]

-

-

-

DR TLMLMT [Mbits]

-

-

-


daily note

[Maintenance Note (Matsuzaki), for CO/CP/RT Tohban]
NFS server for /home/flare/all directories are not available at 10:00-11:30.
The XRT full sun image referred from the pointing tool cannot be
accessed in this period. Please use the other images (SOHO etc...).
No other effects are expected for this activity.

[Maintenance Note (Yamagishi), for CO/CP/RT Tohban]
Some software packages will be updated with command 'yum update' on machines
'cpws', 'xrtco', and 'eisco' at 15:00-18:00(JST) on Tuesday the 9th of December.
When that's over, the machine 'aocswin' will be rebooted.
Please report if you have any trouble.

XRT:

- XBP evolution study (coordinated with EIS)
- XBP brightness study (latitude dependence) - two fixed pointings
- QS dynamics (burst mode obesrvations)
- Disk coronal hole observations (coordinated with EIS)
- Hot thermal/non-thermal studies
- Testing a flare response program in QT mode

Note: XRT is conducting Remote Operations from December 5th through December 12th, 2008. Please continue using Yugma for the Morning Operations Meetings, and dial in to the telecon phone number (see telephone-number card in Hinode morning-meeting room). To contact the XRT Chief Observer, please call +1-617-496-1156 (US phone number).

SOT: HOP 95, magnetic fields of the Network Bright Points (NBP) (SOT only, 2 hours)
Engineering test (2 hours)

EIS:
XBP observations
CH activity study
QS dynamics


[Maintenance Note (Morita), for CO/CP/RT Tohban]
The following machines will be backed up of systems Tuesday,December 9th.
The machines will be rebooted.
Data distributors for real time operation are not related to this maintenance.
Please report if you have any trouble.
-Data distributor used for transfer of commands ; at 9:30-12:00
-ISACS-DOC system ; at 12:00-15:00(The diagnosis service stops.)



CHIEF
TOHBAN

SUB
TOHBAN

CP

SOT CO

XRT CO

EIS CO

EVENING

Arikawa

Okamoto

Cruz

Ishibashi (RCO)

NRL RCO/Brooks

MORNING


operation files

pointing file : none